# CASE – Condition-Aware Sentence Embeddings for Conditional Semantic Textual Similarity Measurement

### v3

@@ Abstract
The meaning conveyed by a sentence often depends on the context in which it appears. Despite the progress of sentence embedding methods, it remains unclear as how to best modify a sentence embedding conditioned on its context. To address this problem, we propose Condition-Aware Sentence Embeddings (CASE), an efficient and accurate method to create an embedding for a sentence under a given condition. First, CASE creates an embedding for the condition using an Large Language Model (LLM), where the sentence influences the attention scores computed for the tokens in the condition during pooling. Next, a supervised non-linear projection is learnt to reduce the dimensionality of the LLM-based text embeddings. We show that CASE significantly outperforms previously proposed Conditional Semantic Textual Similarity (C-STS) methods on an existing standard benchmark dataset. We find that subtracting the condition embedding will consistently improve the C-STS performance of LLM-based text embeddings. Moreover, we propose a supervised dimensionality reduction method that not only reduces the dimensionality of the LLM-based embeddings, but also significantly improves their performance.
@@
Gaifan Zhang1  Yi Zhou2  Danushka Bollegala1 1 University of Liverpool  2 Cardiff University sggzhan8@liverpool.ac.uk, zhouy131@cardiff.ac.uk, danushka@liverpool.ac.uk
@@
Representing the meaning of a given sentence using a vector embedding is a fundamental task required by many Natural Language Processing (NLP) applications Conneau et al. (2017); Reimers and Gurevych (2019); Gao et al. (2021); Xu et al. (2023); Chen et al. (2023). Sentence embeddings are used to measure the Semantic Textual Similarity (STS) between two sentences Agirre et al. (2012, 2015, 2016).
@@
Despite its importance, measuring STS between two sentences is a non-trivial task for humans, which is conditioned on what is being compared between the two sentences. For example, for the two sentences shown in Figure 1, human annotators would assign different similarity ratings, depending on what they are asked to compare (i.e., given conditions 1 and 2). Existing STS benchmarks do not specify the conditions under which two sentences must be compared for their semantic similarity. To address this limitation, Deshpande et al. (2023) proposed the C-STS task and a dataset, where the similarity between two sentences $s_{1}$ and $s_{2}$ is measured under two different conditions $c_{\rm low}$ and $c_{\rm high}$, which focus on different aspects of semantics in the two sentences, resulting in different similarity ratings. Many real-world applications can be seen as C-STS tasks such as ranking a set of documents retrieved for the same query in Information Retrieval (IRManning et al. (2008), comparing two answers for the same question in Question Answering (QARisch et al. (2021), or measuring the strength of a semantic relation between two entities in Knowledge Graph Completion (KGCYoo et al. (2024); Lin et al. (2024).
@@
Figure 1: The two conditions focus on different information described in the two sentences. Human annotators rate the two sentences 1–5, indicating a high-level (5) of semantic textual similarity under condition 2 than condition 1 (1). Our proposed condition-aware sentence embedding (CASE) method reports similarity scores that are well-aligned with the human similarity ratings.
We propose CASE, a method for learning Condition-Aware Sentence Embeddings, for a given input sentence considering another sentence. Specifically, CASE creates an embedding for the condition sentence using an LLM with a prompt that includes the target sentence (e.g. Sentences 1/2 in Figure 1), where the latter is not encoded explicitly in the embedding but influences the attention scores used during token pooling. Compared to Masked Language Models that have been used extensively in prior work on C-STS, LLMs contain billions of parameters and are typically trained on much larger datasets for a longer period of time. Consequently, LLM-based embedding models have consistently ranked at the top in leaderboards evaluating text embedding models Muennighoff et al. (2022). Therefore, by leveraging LLM-based embeddings, CASE is able to benefit from the rich world knowledge contained in LLMs. However, the optimal method to use LLMs for C-STS remains elusive as reported by Lin et al. (2024) who showed that decoder-only LLMs often underperform MLM-based embeddings in C-STS benchmarks. Interestingly, we find that encoding the condition given the sentence often outperforms the reverse setting in C-STS benchmarks. Moreover, we find that subtracting the embedding of the condition in a post-processing step further improves the performance across multiple LLM-based embedding models.
One disadvantage of using LLM-based embeddings compared to their MLM-based counterparts is the high dimensionality of the LLM-based embeddings. For example, embeddings produced by state-of-the-art (SoTA) LLM encoders such as NV-embed-v2 Lee et al. (2024) are 4096 dimensional, whereas MLM embeddings such as RoBERTa-base Liu et al. (2019) SimCSE Gao et al. (2021) embeddings are 768 dimensional. High dimensional embeddings are problematic both due to their high storage requirements and the high inner-product computation cost. Moreover, the embeddings obtained from LLMs are not necessarily aligned to the C-STS task because they are not fine-tuned on such tasks. To address these issue, we propose a supervised dimensionality reduction method that accurately projects LLM-based embeddings to low dimensional vector spaces, while also improving their performance in the C-STS task. We find that although popular unsupervised dimensionality reduction methods such as Principal Component Analysis (PCA) and Independent Component Analysis (ICA) are effective at reducing the dimensionality of the embedding space, they underperform our proposed supervised projection learning method. Further investigations into individual examples show that CASE increases the similarity between a sentence and the information emphasized by a given condition, while decreasing the same for information irrelevant to the condition as expected. We have submitted the source code and pre-processed data for reproducing our findings for anonymous reviewing, which will later be publicly released upon paper acceptance.
@@
Conditional Semantic Textual Similarity:
Deshpande et al. (2023) proposed the C-STS task for measuring the similarity between two sentences under a given condition. They created a human-annotated dataset containing 18,908 instances where the semantic similarity between two sentences $s_{1}$ and $s_{2}$ is rated under two conditions $c_{\rm high}$ and $c_{\rm low}$ resulting in respectively high vs. low similarity between the two sentences. Moreover, they proposed cross-, bi- and tri-encoder baselines. Given a triplet $(s_{1},s_{2},c)$ a cross-encoder considers all interactions among tokens in $s_{1}$, $s_{2}$ and $c$. Although the cross-encoder setting benefits from having access to both $s_{1}$ and $s_{2}$ at the same time, it is computationally costly due to the large number of token interactions required for longer sentences and conditions. Moreover, it does not pre-compute conditional embeddings for the individual sentences, which is problematic when computing all pairwise similarities between the sentences in a large set.
Bi-encoders overcome the limitations of cross-encoders by creating a condition-aware embedding for each sentence, and then compute C-STS using an efficient operation such as the inner-product. Tri-encoders separately encode $s_{1}$, $s_{2}$ and $c$, and then apply some late interactions between the condition’s and each sentence’s embeddings to compute C-STS. Despite the computational benefits gained by pre-encoding sentences and conditions separately, late interaction mechanisms for tri-encoders remain complex and under-perform in C-STS benchmarks.
@@
Tu et al. (2024) found that the C-STS dataset created by Deshpande et al. (2023) contain ill-defined conditions and annotation errors, resulting in a significant discrepancy among the annotators for over half of the dataset. To address this, Tu et al. (2024) re-annotated the validation split from the original C-STS dataset. Moreover, they proposed a method to solve C-STS by first converting each condition into a question, and then using GPT-3.5 to extract the corresponding answers from the two sentences. Finally, C-STS is estimated as the cosine similarity between the SimCSE embeddings for the two answers. Although this QA formulation consistently improves the performance of all baseline encoders, it depends on multiple decoupled components such as converting conditions into questions, requiring a decoder LLM to extract the answer from each question, and using a separate encoder to compare the extracted answers, which increases the possibility of error propagation across components.
Yoo et al. (2024) proposed Hyper-CL, a contrastively learnt hypernetwork Ha et al. (2017) to selectively project the embeddings of $s_{1}$ and $s_{2}$ according to $c$. Hyper-CL follows a tri-encoder setting where $s_{1}$, $s_{2}$ and $c$ are first encoded separately using a sentence encoder. Next, a hypernetwork is trained to produce a linear transformation matrix conditioned on $c$. Finally, the embeddings of $s_{1}$ and $s_{2}$ are projected using this transformation matrix and their inner-product is computed. Hyper-CL improves the performance of tri-encoder models, but still under-performs bi-encoders for C-STS. Moreover, hypernetworks introduce an external parameter set that is three times larger than the SimCSE model used to encode each sentence, resulting in an excessively large memory space.
Lin et al. (2024) proposed a tri-encoder-based C-STS method where they used routers and heavy-light attention Ainslie et al. (2023) to select the relevant tokens to a given condition. Specifically, they used the query vector of the [CLS] token of the condition to compute attention scores for the tokens in the sentence, which are subsequently used to compute a sentence embedding. Their method outperforms Hyper-CL for the tri-encoder setting. Liu et al. (2025) proposed a conditional contrastive learning method for C-STS, introducing weighted contrastive losses with a sample augmentation strategy. Although it improved performance for both bi- and tri-encoders, the former outperforms the latter.
Li et al. (2024a) proposed a cross-encoder approach which predicts C-STS scores, without creating conditional embeddings. Prior work on C-STS has shown cross-encoders to perform poorly despite having access to both sentences simultaneously. Li et al. (2024a) showed that this is due to the irrelevant information in the two sentences to the condition and proposed a token re-weighting strategy, inspired by object detection in computer vision Shi et al. (2023); Jaegle et al. (2021). Concretely, they compute two cross-attention matrices between $(s_{1},s_{2})$ and $c$, which are subsequently used to compute the correlations for the sentence or condition tokens. Although their method improves the performance of cross-encoder-based C-STS measurement, it still underperforms bi-encoders.
Figure 2: Overview of CASE. An LLM is prompted with $I(s)$ to obtain two separate embeddings $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ for the condition $c$ for two sentences $s_{1}$ and $s_{2}$. The unconditional embedding $f(c;I(\emptyset))$ is then computed using the prompt $I(\emptyset)$ and subtracted from each of those embeddings. Finally, the embeddings are projected to a lower-dimensional space using a learnt projection $g$ and their cosine similarity is computed.
Text Embeddings from LLMs:
MLMs use a bi-directional attention Devlin et al. (2018), where the information from tokens appearing in positions both before and after the current token are used to predict the embedding for the current target token in the sequence that must be encoded. In contrast, decoder-only LLMs are trained with a causal attention mask Vaswani et al. (2017), which prevents information leakage from future tokens by allowing the decoder to attend only to previous positions during auto-regressive text generation. Although this makes sense for decoders, it is sub-optimal when using LLMs for encoding a given sequence as evident from the poor performance of GPT models Lin et al. (2024) compared to similar-sized BERT Devlin et al. (2018) or T5 Chung et al. (2024) models on various natural language understanding benchmarks Wang et al. (2019).
BehnamGhader et al. (2024) proposed a post-processing method for obtaining text embeddings from decoder-only LLMs following three steps: (a) enable bidirectional attention to overcome the restrictions due to causal attention, (b) train the model to predict the masked next token using bi-directional attention, and (c) use unsupervised contrastive learning to compute better sequence representations. Finally, a pooling method is applied on the token embedding sequence to create a fixed-dimensional embedding for the input text such as the embedding of the last token in the sequence (i.e. last token pooling) Meng et al. (2024); Li et al. (2023) or the average over all token embeddings (i.e. mean pooling) Wang et al. (2024). Moreover, Lee et al. (2024) proposed a latent attention layer where they compute the cross attention between the last hidden layer from the decoder and a trainable latent array to compute a weighted pooling mechanism.
@@
An overview of our proposed method is shown in Figure 2, which consists of two-steps. In the first step (§ 3.1), we create two separate embeddings for the condition considering each of the two sentences, one at a time, in the instruction prompt shown to an LLM. Note that it is the condition that is being encoded and the tokens in the sentence (similar to all other tokens in the instruction) are simply modifying the attention scores computed for the tokens in the condition. Intuitively, it can be seen as each sentence filling some missing information asked in the condition.
Embeddings obtained from LLMs are typically high dimensional and not necessarily fine-tuned for the C-STS task. Therefore, in the second step (§ 3.2), we learn a supervised projection using the training split from the C-STS dataset to project each of the two condition embeddings to a lower-dimensional vector space. Finally, the C-STS between two sentences is computed as the cosine similarity between the corresponding projected embeddings.
@@
Given an LLM-based encoder, $f$, we create a $d$-dimensional embedding $f(c;I(s))\in\mathbb{R}^{d}$ for a condition $c$, given the sentence $s$. Here, $I$ is an instruction template that takes $c$ as an argument. We use the following prompt template as $I(s)$Retrieve semantically similar texts to the Condition, given the Sentence : [SENTENCE], where we substitute $s$ in the placeholder [SENTENCE]. Next, we provide $c$ as the input text to be encoded by the LLM following the instruction $I(s)$. Finally, the token embeddings of $c$ are aggregated according to one of the pooling methods described previously to create $f(c;I(s))$. Although our focus here is to create condition-aware embeddings, it is noteworthy that we can also obtain unconditional embeddings for a sentence by dropping the condition in the above prompt. Specifically, we use the prompt $I(\emptyset)$ Retrieve semantically similar texts to a given Sentence for this purpose, and denote this unconditional embedding of $c$ by $f(c;I(\emptyset))$. As we see later in our experiments, by subtracting $f(c;I(\emptyset))$ from $f(c;I(s))$, we can reduce the effect of tokens in the condition that are irrelevant to the sentence, thereby consistently improving accuracy of the condition-aware embeddings. This first step is fully unsupervised and a zero-shot prompt template is used as $I$.
Recall that both $s$ and $c$ are text strings, and it is possible to swap the sentence and condition in the above formulation to obtain an embedding $f(s;I(c))$ for the sentence, given the condition. However, as shown later in our experiments, comparing the embedding for $c$ created under $s_{1}$ and $s_{2}$ results in better performance on the C-STS benchmark for all LLM encoders. This is because $s_{1}$ and $s_{2}$ contain many irrelevant information to $c$, which will affect the cosine similarity computed between $f(s_{1};I(c))$ and $f(s_{2};I(c))$. On the other hand, the cosine similarity between $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ is a more accurate estimate of the C-STS between $s_{1}$ and $s_{2}$ under $c$ because it is purely based on the meaning alterations to $c$ under $s_{1}$ and $s_{2}$ separately.
@@
The LLM-based embeddings computed in § 3.1 has two main drawbacks. First, relative to MLMs-based sentence embeddings, LLMs produce much higher dimensional embeddings, which can be problematic due to their memory requirements (especially when operating on a limited GPU memory) and the computational cost involved in inner-product computations. In tasks such as dense retrieval, we must compare millions of documents against a query to find the nearest neighbours under strict latency requirements, and low-dimensional embeddings are preferable. Second, although LLMs are typically trained on massive text collections and instruction-tuned for diverse tasks Muennighoff et al. (2022), their performance on C-STS tasks have been poor Lin et al. (2024). As seen from our condition-aware prompt template, an LLM must be able to separately handle a variable condition statement and a fixed instruction. This setup is different from most tasks on which LLMs are typically trained on, where the instruction remains fixed across all inputs. Therefore, it is important to perform a supervised fine-tuning step to LLM embeddings before they are used for C-STS.
To address the above-mentioned drawbacks, we propose a supervised projection learning method. Specifically, we freeze the model parameters of the LLM and use an Multi-layer Perception (MLP) layer that takes $f(c;I(s))$ as the input and returns a $k$-dimensional ($k\leq d)$ embedding $g(f(c;I(s);\theta)$, where $\theta$ denotes the learnable parameters of the MLP. Finally, we define ${\rm CASE}(s,c)$, as the projection of the offset between the conditional and the unconditional embeddings of $c$ under $s$, given by (1).
${\rm CASE}(s,c)=g(f(c;I(s_{1}))-f(c;I(\emptyset));\theta)$ (1)
We use the human similarity ratings $r$ in the C-STS train instances $\mathcal{D}$ to learn $\theta$. Specifically, we minimise the squared error between the human ratings and the cosine similarity computed using the corresponding CASE as given by (2).
@@
Here, $\cos$ denotes the cosine similarity between the projected embeddings. We use Adam optimiser Kingma and Ba (2014) to find the optimal $\theta$ that minimises the loss given by (2). Recall that only the MLP parameters are updated during this projection learning step, while keeping the parameters of the LLM fixed, which makes it extremely efficient. For example, it takes less than 5 minutes to learn this projection even for the largest (4096 dimensional) embedding spaces using the train split of the C-STS dataset. Using the learnt projection, we compute the C-STS between $s_{1}$ and $s_{2}$ under $c$ as the cosine similarity between ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$.
@@
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we apply six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional NV), SFR-Embedding-Mistral (4096 dimensional SFR), gte-Qwen2-7B-instruct (3584 dimensional GTE) and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional E5), sup-simcse-roberta-large (1024 dimensional SimCSE_large), and sup-simcse-bert-base-uncased (768 dimensional SimCSE_base). Further details provided in Appendix A.
@@
We evaluate model performance on different pooling methods, prompt settings, and sentence constructions. Moreover, we evaluate the dimensionality reduction methods under supervised and unsupervised settings to learn the projection for CASE (PCA, ICA, and linear and non-linear MLPs). Both the linear and non-linear MLPs are Siamese bi-encoders with weight-sharing. As explained in § 3.2, they take $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ as the input embeddings, and return the projected embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$ as the outputs, which can be compared using a similarity metric such as cosine.
Linear MLP performs a linear transformation:
$\mathbf{z}=\operatorname{Dropout}\left(\text{\bf W}\bm{e}\right)$ (3)
@@
where $\text{\bf W}\in\mathbb{R}^{d^{\prime}\times d}$ is the learned projection matrix. A dropout layer is applied to reduce any overfitting Hinton et al. (2012).
Our non-linear MLP is a two-layer MLP:

### v4

@@ Abstract
The meaning conveyed by a sentence often depends on the context in which it appears. Despite the progress of sentence embedding methods, it remains unclear as how to best modify a sentence embedding conditioned on its context. To address this problem, we propose Condition-Aware Sentence Embeddings (CASE), an efficient and accurate method to create an embedding for a sentence under a given condition. First, CASE creates an embedding for the condition using a Large Language Model (LLM) encoder, where the sentence influences the attention scores computed for the tokens in the condition during pooling. Next, a supervised method is learnt to align the LLM-based text embeddings with the Conditional Semantic Textual Similarity (C-STS) task. We find that subtracting the condition embedding consistently improves the C-STS performance of LLM-based text embeddings by improving the isotropy of the embedding space. Moreover, our supervised projection method significantly improves the performance of LLM-based embeddings despite requiring a small number of embedding dimensions.111code: https://github.com/LivNLP/CASE
@@
Gaifan Zhang1   Yi Zhou2   Danushka Bollegala1,3 1 University of Liverpool  2 Cardiff University  3 Amazon {sggzhan8,danushka}@liverpool.ac.uk, zhouy131@cardiff.ac.uk
@@
Measuring Semantic Textual Similarity (STS) between sentences is a fundamental task in NLP Majumder et al. (2016). It is also important in training the sentence encoders Reimers and Gurevych (2019). Recent approaches, such as contrastive learning, use semantic similarity as an objective to improve the quality of sentence embeddings Gao et al. (2021). However, measuring sentence similarity is a complex task, which depends on the aspects being considered in the sentences being compared. To address this problem, Deshpande et al. (2023) proposed the C-STS task along with a human-annotated dataset. They designed the C-STS task and the dataset by assigning different conditions (semantic aspects) to each pair of sentences ($s_{1}$, $s_{2}$), a condition of high similarity $c_{\rm high}$ and a condition of low similarity $c_{\rm low}$, resulting in different similarity ratings for the same pair of sentences. This reduces subjectivity and ambiguity in the measurement of similarity between two sentences. As shown in Figure 1, human annotators are required to assign different similarity scores under different conditions, rather than a single score as in the traditional STS task. Many real-world applications can be seen as C-STS tasks such as ranking a set of documents retrieved for the same query in Information Retrieval (IR) Manning et al. (2008), comparing two answers to the same question in Question Answering (QA) Risch et al. (2021), or measuring the strength of a semantic relation between two entities in Knowledge Graph Completion (KGC) Yoo et al. (2024); Lin et al. (2024).
@@
Figure 1: The two conditions focus on different information described in the two sentences. Human annotators rate the two sentences 1–5, indicating a high-level (5) of semantic textual similarity under $c_{\rm high}$ than $c_{\rm low}$ (1). Our proposed condition-aware sentence embedding (CASE) method reports similarity scores that are well-aligned with the human similarity ratings.
Figure 2: Overview of CASE. An LLM is prompted with $I(s)$ to obtain two separate embeddings $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ for the same condition $c$ for the two sentences $s_{1}$ and $s_{2}$. The unconditional embedding $f(c;I(\emptyset))$ is then computed using the prompt $I(\emptyset)$ and subtracted from each of those conditional embeddings. Finally, the embeddings are projected to a lower-dimensional space using a supervised FFN and their cosine similarity is computed as the C-STS rating.
We propose CASE (Condition-Aware Sentence Embeddings), an efficient two-step method to combine LLM-based sentence encoders and a lightweight supervised projection, to create the embeddings required for the C-STS task (see Figure 2). In the first step, we prompt the LLM encoders to create the initial embeddings, with an instruction based on conditional semantic similarity to encode the condition given a sentence. Here, the sentence is not encoded explicitly but instead influences the attention scores during token pooling. Compared to Masked Language Model (MLM)-based encoders that have been used in prior work on C-STS, LLM-based encoders can be used as accurate embedding models by direct prompting Tao et al. (2025). They benefit from large-scale datasets and complex architectures with billions of parameters, excelling for various downstream tasks, as demonstrated by the performance on the MTEB leaderboard Muennighoff et al. (2023). However, the best way to use LLMs for C-STS remains elusive, as reported by Lin et al. (2024), who showed that decoder-only LLMs often underperform MLM-based embeddings in C-STS benchmarks. Interestingly, we find that our prompt structure “retrieve [condition], given [sentence]” counter-intuitively performs better than “retrieve [sentence], given [condition]”. Furthermore, a post-processing step of subtracting the embedding of the condition significantly improves the performance and isotropy of the embedding space (as shown in § 4).
Previous C-STS methods fine-tune sentence encoders Li et al. (2024); Tu et al. (2024); Yoo et al. (2024); Lin et al. (2024), which is computationally expensive, particularly for LLM encoders due to their significantly larger number of parameters (>1B), even with parameter-efficient fine-tuning methods. Moreover, LLM encoders produce high-dimensional embeddings (e.g. 4096 for NV-embed-v2 Lee et al. (2024)) compared to MLM encoders (e.g. 768 for SimCSE Gao et al. (2021)), posing challenges when storing and computing inner-products between embeddings. To address these issues, the second step of CASE trains a lightweight supervised Feed Forward Network (FFN) (<5M) in a bi-encoder setting to fine-tune the embeddings in a lower-dimensional space. Qualitative analysis shows that CASE increases the similarity between a sentence and the information emphasised by a given condition, while decreasing the same for information irrelevant to the condition, as expected.
Our CASE method makes the following contributions:
  • It uses the superior semantic capabilities of LLM-based embedding models.
  • It avoids fine-tuning or re-training existing encoders by using a lightweight FFN.
[…]
@@
Deshpande et al. (2023) proposed the C-STS task and created a dataset containing 18,908 instances where the semantic similarity between two sentences $s_{1}$ and $s_{2}$ is rated under two conditions $c_{\rm high}$ and $c_{\rm low}$ resulting in, respectively, high vs. low similarity between the two sentences. Moreover, they proposed cross-, bi- and tri-encoder baselines. Cross-encoders consider interactions between all tokens in $s_{1}$, $s_{2}$ and $c$, which are computationally expensive for large-scale comparison due to a lack of pre-computed conditional embeddings. Tri-encoders separately encode $s_{1}$, $s_{2}$ and $c$, and then apply late interactions between the condition’s and each sentence’s embeddings, which are also complex and report suboptimal performance. Bi-encoders overcome those limitations by creating a condition-aware embedding for each sentence and computing C-STS efficiently as their inner-product.
@@
Tu et al. (2024) found multiple issues in the dataset created by Deshpande et al. (2023), such as poorly defined conditions and inconsistent similarity ratings for more than half of the dataset. To address this, Tu et al. (2024) re-annotated the validation split from the original C-STS dataset. Moreover, they proposed a QA-based approach for measuring C-STS by first converting conditions into questions, and then using GPT-3.5 to extract the corresponding answers, which are the compared using their embeddings. This QA formulation depends on multiple decoupled components, such as converting conditions into questions, requiring a decoder LLM to extract answers, and using a separate encoder to generate embeddings, which increases the possibility of propagation of errors between components.
Zhang et al. (2025) further investigated the dataset created by Deshpande et al. (2023) and found issues such as problematic condition statements and inaccurate human annotations. They proposed an LLM-based two-step method to improve the dataset by first refining the condition statements and then re-annotating the ratings, creating a larger scale (14176 instances) dataset of better quality, which we use in our experiments as training data.
Alternative architectures to the tri-encoder and cross-encoder have been explored to capture conditional semantics. Yoo et al. (2024) proposed Hyper-CL, a tri-encoder using contrastively learnt hypernetwork Ha et al. (2017) to selectively project the embeddings of $s_{1}$ and $s_{2}$ according to $c$. Hypernetworks introduce an external parameter set that is three times larger than that of the SimCSE model used to encode each sentence, requiring in an excessively large memory space, which is problematic when processing large sets of sentences. Lin et al. (2024) proposed a tri-encoder-based C-STS method where they used routers and heavy-light attention Ainslie et al. (2023) to select the relevant tokens to a given condition. Liu et al. (2025) proposed a conditional contrastive learning method for C-STS, introducing a weighted contrastive loss with a sample augmentation strategy. Li et al. (2024) proposed a cross-encoder approach, which predicts C-STS scores, without creating conditional embeddings. They used a token re-weighting strategy by computing two cross-attention matrices between $(s_{1},s_{2})$ and $c$, which are subsequently used to compute the correlations for the sentence or condition tokens. Although the aforementioned methods improve the performance of cross-encoder and tri -encoder-based C-STS measurement, they still underperform bi-encoders.
Li and Li (2024) proposed BeLLM, which uses backward dependency to enhance LLMs by learning embeddings from uni- to bi-directional attention layers. They fine-tuned the LLaMA2-7B model with their method on the C-STS dataset, outperforming previously proposed fine-tuned models. In particular, they designed a prompt template for this task: “Given the context [condition], summarise the sentence [sentence] in one word:”. They then extracted embeddings from the hidden states of the generated output and used these embeddings to measure sentence similarity. Their results showed that larger model sizes and backward dependencies contribute to the fine-tuning effects. In contrast to our CASE, which uses LLM encoders, they use decoder-only LLMs. Moreover, we do not require token generation and therefore are not restricted to single-word summarisation. Additionally, we find that the reversed structure (“given [sentence], retrieve [condition]”) performs better.
Yamada and Zhang (2025) proposed an unsupervised conditional text embedding method that uses a causal LLM. Specifically, they use the input prompt “Express this text [sentence] in one word in terms of [condition],” and compute the cosine similarity of the last token’s hidden state as the C-STS score. They further validated that instruction-tuned models outperform their non-instruction-tuned counterparts. Similarly, our CASE leverages prompt-based LLM encoders to generate conditional embeddings, yet achieves performance comparable to the best supervised bi-encoder settings in a low-dimensional space.
@@
An overview of our proposed method is shown in Figure 2, which consists of two-steps. In the first step (§ 3.1), we create two separate embeddings for the condition considering each of the two sentences, one at a time, in the instruction prompt shown to an LLM. Note that it is the condition that is being encoded and the tokens in the sentence (similar to all other tokens in the instruction) are simply modifying the attention scores computed for the tokens in the condition. Intuitively, it can be seen as each sentence filling some missing information required by the condition. However, note that the embeddings obtained from LLMs are not necessarily aligned with the C-STS task. Therefore, in the second step (§ 3.2), we learn a projection layer using the training split from the C-STS dataset. Finally, the C-STS between two sentences is computed as the cosine similarity between the corresponding projected embeddings.
@@
Given an LLM-based encoder, $f$, we create a $d$-dimensional embedding that captures the semantic relationship between a sentence $s$ and a condition $c$. There are two ways to formulate the input:
  1. 1.
    Encode the condition given the sentence, denoted by $f(c;I(s))$. Here, $I$ is an instruction template that takes $c$ as an argument. We use the following prompt template as I(s)I(s)Retrieve semantically similar texts to a given Condition, given the Sentence : [s], where we substitute $s$ in the placeholder [s]. Next, we provide $c$ as the input text to be encoded by the LLM following the instruction $I(s)$. Finally, the token embeddings of $c$ are aggregated according to one of the pooling methods to create $f(c;I(s))$ (different pooling methods are discussed in Appendix D).
  2. 2.
    Encode the sentence given the condition, denoted by $f(s;I(c))$. Here we swap the sentence and the condition in the above formulation, recalling that both $s$ and $c$ are text strings. In particular, as shown later in our experiments (§ 4.2), comparing the embedding for $c$ created given $s_{1}$ and $s_{2}$ results in better performance on the C-STS benchmark for all LLM encoders. For this reason, we keep the best-performing $f(c;I(s))$ in Figure 2 and our method for subsequent experiments. This is because $s_{1}$ and $s_{2}$ will also contain information irrelevant to $c$, which will affect the cosine similarity computed between $f(s_{1};I(c))$ and $f(s_{2};I(c))$. On the other hand, the cosine similarity between $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ is a more accurate estimate of C-STS between $s_{1}$ and $s_{2}$ under $c$ because it is purely based on contextualised representation shifts of $c$ conditioned on $s_{1}$ and $s_{2}$ separately.
Condition statements contain their own intrinsic semantics (e.g., the phrase colour of dress), which is irrelevant when distinguishing specific semantic values based on sentences (e.g., orange vs. grey). To address this problem, we introduce a post-processing step of subtracting the embedding of the condition to reduce the effect of tokens in the condition that are irrelevant to the sentence, thereby consistently improving the accuracy of the condition-aware embeddings. Specifically, we use the prompt $I(\emptyset)$ Retrieve semantically similar texts to obtain the embedding of a condition $c$, and denote this by $f(c;I(\emptyset))$. As we see later in our experiments, by subtracting $f(c;I(\emptyset))$ from $f(c;I(s))$, we improve the performance on the C-STS task and the isotropy of embeddings. This first step is fully unsupervised and a zero-shot prompt template is used as $I$.
@@
The LLM embeddings computed in § 3.1 has two main drawbacks. First, although LLMs are typically trained on massive text collections and instruction-tuned for diverse tasks Muennighoff et al. (2023), their performance on C-STS tasks have been poor Lin et al. (2024). As seen from our condition-aware prompt template, an LLM must be able to separately handle a variable condition statement and a fixed instruction. This setup is different from most tasks on which LLMs are typically trained on, where the instruction remains fixed across all inputs. Second, relative to MLMs-based sentence embeddings, LLMs produce much higher dimensional embeddings, which can be problematic due to their memory requirements (especially when operating on a limited GPU memory) and the computational cost involved in inner-product computations. In tasks such as dense retrieval, we must compare millions of documents against a query to find the nearest neighbours under strict latency requirements, and low-dimensional embeddings are preferable.
To address the above-mentioned drawbacks, we propose a supervised projection learning method. Specifically, we freeze the model parameters of the LLM and use a FFN layer that takes $f(c;I(s))$ as the input and returns a $k$-dimensional ($k\leq d)$ embedding $g(f(c;I(s);\theta)$, where $\theta$ denotes the learnable parameters of the FFN. Finally, we define ${\rm CASE}(s,c)$ as the projection of the offset between the conditional and the unconditional embeddings of $c$ under $s$, given by,
${\rm CASE}(s,c)=g(f(c;I(s))-f(c;I(\emptyset));\theta)$ (1)
We use the human-annotated similarity ratings $r$ in the C-STS train instances $\mathcal{D}$ to learn $\theta$. Specifically, we minimise the squared error between the human ratings and the cosine similarity computed using the corresponding CASE as given by (2).
@@
Here, $\cos$ denotes the cosine similarity between the projected embeddings. We use Adam optimiser Kingma and Ba (2014) to find the optimal $\theta$ that minimises the loss given by (2). Recall that only the FFN parameters are updated during this projection learning step, while keeping the parameters of the LLM fixed, which makes it extremely efficient. Using the learnt projection, we compute the C-STS between $s_{1}$ and $s_{2}$ under $c$ as the cosine similarity between the embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$.
@@
To evaluate the effectiveness of the LLM-based and MLM-based sentence embeddings as described in § 3, we use six sentence encoders, out of which three are LLM-based: NV-Embed-v2 (4096 dimensional NV), SFR-Embedding-Mistral (4096 dimensional SFR), gte-Qwen2-7B-instruct (3584 dimensional GTE) and three are MLM-based: Multilingual-E5-large-instruct (1024 dimensional E5), sup-simcse-roberta-large (1024 dimensional SimCSE_large), and sup-simcse-bert-base-uncased (768 dimensional SimCSE_base). Further details provided in Appendix A.
Model sent/cond? Spear.
NV sent - c 16.98
sent 22.07
cond - c 37.61
cond 33.85
SFR sent - c 19.54
sent 11.89
cond - c 20.44
cond 18.32
GTE sent - c 7.16
sent 7.16
cond - c 20.40
cond 16.58
E5 sent - c 11.08
sent 3.77
cond - c 15.37
cond 6.07
SimCSE_large $\operatorname{CONC}(s+c)$ 5.59
$\operatorname{CONC}(c+s)$ 4.00
$\operatorname{CONC}(s+c)$ - c 8.32
$\operatorname{\textbf{CONC}}\bm{(c+s)}$ - c 8.58
SimCSE_base $\operatorname{CONC}(s+c)$ 4.37
$\operatorname{CONC}(c+s)$ 1.25
$\operatorname{\textbf{CONC}}\bm{(s+c)}$ - c 7.05
$\operatorname{CONC}(c+s)$ - c 6.00
Table 1: Spearman scores for different sentence embedding models and encoding settings.
@@
We evaluate model performance on different pooling methods, prompt settings, and sentence constructions. Moreover, we evaluate the supervised methods of linear and non-linear FFNs. Both the linear and non-linear FFNs are Siamese bi-encoders with weight-sharing. As explained in § 3.2, they take $f(c;I(s_{1}))$ and $f(c;I(s_{2}))$ as the input embeddings, and return the projected embeddings ${\rm CASE}(s_{1},c)$ and ${\rm CASE}(s_{2},c)$ as the outputs, which can be compared using a similarity metric such as the cosine similarity.
Linear FFN performs a linear transformation:
$\mathbf{z}=\text{\bf W}\bm{e}$ (3)
@@
where $\text{\bf W}\in\mathbb{R}^{d^{\prime}\times d}$ is the learned projection matrix.
Our non-linear FFN is a one-layer FFN with a activation function LeakyReLU: